1) Please calculate the number of moles that are in 10.0 mL of water (H2O).

2) please calculate the mixed volume (Vmix) when 0.300 mol of ethanol and 0.300 mol of water are mixed.

a) 74.3mL

b)22.3mL

c) 22.3mL

d) 37.2mL

3) Please calculate the mole fraction of ethanol and the mole fraction of water when

0.300 mol of ethanol and 0.200 mol of water are mixed.

mole fraction of ethanol = 0.150 mole fraction of water = 0.150

mole fraction of ethanol = 0.600       mole fraction of water = 0.400

mole fraction of ethanol = 0.500       mole fraction of water = 0.500

mole fraction of ethanol = 0.300       mole fraction of water = 0.300
Answer #1

1) The moles of water are calculated:

n water = 10 mL * (1 g / 1 mL) * (1 mol / 18 g) = 0.56 mol

2) The volume of water and ethanol and the mixture are calculated:

V water = 0.3 mol water * (18 g / 1 mol) * (1 mL / 1 g) = 5.2 mL

V ethanol = 0.3 mol * (46.07 g / 1 mol) * (1 mL / 0.789 g) = 17.1 mL

V mix = 5.2 + 17.1 = 22.3 mL

3) The molar fraction is calculated:

X water = n water / n total = 0.2 / 0.5 = 0.400

X Ethanol = 1 - 0.4 = 0.600

If you liked the answer, please rate it positively, you would help me a lot, thanks.